Find the linear equation that passes through the point (-8,1) and has an undefined slope.
Write the equation of the line parallel to x=8 that passes through the point (-3,0).
Find the slope of the line passing through the points (-4,2) and (1,-5).

Answer by texttutoring(324)

Undefined slope means that the slope is dividing by zero. So there is no 'run'. It is a straight up and down, vertical line. If it passes through (-8,1) then the line must be x = -8.

A line parallel to x = 8 would also have an undefined slope, meaning it is a vertical line. If it passes through the point (-3,0) then it must be x=-3. It's a vertical line at x=-3.

Lastly, if a line passes through (-4,2) and (1,-5). First find the slope:

m = (y2-y1) / (x2-x1)
m = (-5 - 2) / (1 - -4)
m = -7 / (1+4)
m = -7 / 5

Now put this value for m into y=mx+b, and use either point, (-4,2) or (1,-5) to solve for b:

y=mx+b
-5 =(-7/5)(1) + b
-5 + 7/5 = b
b = -18/5

The equation of the line is thus y = (-7/5)x -18/5

Hope this helps! Let me know if you're confused by anything I wrote.